I am having a hard time finishing this problem.
I need to find the equation of the line containing the given pair of points.
(-7,-3) and (-5,-4)
This is what I get so far.
-4-(-3) 1
------- = - -
-5(-7) 2
y-3=m (x-x)
y-3= 1/2 (x-x)
y-3= 1/2 (x-(-7)
y-3= 1/2 - ?? And this is where I keep getting lost I know I am suppose to multiply 1/2 and -7 but I keep getting the wrong answer somehow each time.... I know I am then suppose to subtract 3 from both sides to get y by itself but I am confused by then..I know the final equation should be I believe y= 1/2x- ??

Thank you

Answer by Alan3354(69443)

I need to find the equation of the line containing the given pair of points.
(-7,-3) and (-5,-4)
This is what I get so far.
-4-(-3) 1
------- = - -
-5(-7) 2
Slope m = (-4 - (-3))/(-5 - (-7)) = -1/2, not +1/2
----------------
y+3 = m(x+7)
y+3 = (-1/2)(x+7)
2y + 6 = -x - 7
x + 2y = -13
----------------
Or, use y = mx + b and either point to find b
-3 = (-1/2)*(-7) + b
b = -3 - 7/2
--> y = (-1/2)x - 13/2
----------------------------
----------------------------
y-3=m (x-x)
y-3= 1/2 (x-x)
y-3= 1/2 (x-(-7)
y-3= 1/2 - ?? And this is where I keep getting lost I know I am suppose to multiply 1/2 and -7 but I keep getting the wrong answer somehow each time.... I know I am then suppose to subtract 3 from both sides to get y by itself but I am confused by then..I know the final equation should be I believe y= 1/2x- ??